If the masses of two planets doubles and the distance between them reduces by half, the force between them would be
Answer #1

let

m1= mass of the planet 1st

m2= mass of the second planet

d= distance between the planets

force between two planets is

F = (G m1 m2 )/(d2 )

now the masses of the planets  becomes doubled and distance is reduce to half then

"m1 " becomes "2m1 " and "m2 " becomes "2m2" and "d" becomes "d/2" then force becomes

F' = (G 2m12m2) / (d/2)2 = 4Gm1m2 / ( d2/4)

F' = 16(G m1 m2 )/(d2 ) = 16F

the force between them would be BECOMES 16 TIMES OF INITIAL FORCE